OpenStudy (anonymous):

3(-y + 4) - 12y - 7 show work please

OpenStudy (saifoo.khan):

Try it Austin.

OpenStudy (anonymous):

yes, if your are just simplifying.. distribute

OpenStudy (anonymous):

3(-y+4)-12y-7= -3y+12-12y-7= 9y=-5 divide by 9 to get :-5/9=y